OpenStudy (anonymous):

Find the inverse of f(x) = (5x-4)/(x+4)

OpenStudy (anonymous):

I want to see if I'm on the right track... y = 5x-4/x+4 (y+4)*x = 5y-4 xy + 4x - 4x = 5y - 4 xy = 5y-4-4x

jigglypuff314 (jigglypuff314):

that is correct so far

jigglypuff314 (jigglypuff314):

now get the 5y to the other side and factor out the y

OpenStudy (anonymous):

xy-5y=-4-4x

OpenStudy (anonymous):

5(y-x)=-4-4x

OpenStudy (anonymous):

Oh I factored out the wrong thing

OpenStudy (anonymous):

y(-5+x)=-4-4x

jigglypuff314 (jigglypuff314):

yes then divide (-5+x)

OpenStudy (anonymous):

y = -4-4x/-5+x

jigglypuff314 (jigglypuff314):

that is correct

OpenStudy (anonymous):

wow that is a lot of steps for this one

OpenStudy (anonymous):

could I have taken a shortcut?

jigglypuff314 (jigglypuff314):

I don't see any short cuts...

jigglypuff314 (jigglypuff314):

you could, however, factor out and cancel out some negatives

OpenStudy (anonymous):

can you give me an example?

jigglypuff314 (jigglypuff314):

I got f^-1(x) = 4(x+1) / (5-x)

OpenStudy (anonymous):

That is a nicer solution... mine is (-4-4x)/(-5+x) and it was correct

jigglypuff314 (jigglypuff314):

yes it was

OpenStudy (anonymous):

now I wonder how you got your answer

jigglypuff314 (jigglypuff314):

........ (y+4)*x = 5y-4 xy + 4x = 5y - 4 -xy -xy 4x = 5y - xy -4 +4 +4 4x + 4 = 5y - xy 4x + 4 = y(5-x) y = (4x+4)/(5-x)

OpenStudy (anonymous):

Oh so you subtracted xy first and I subtracted 4x